Consider the following information for Maynor Company, which uses a periodic inventory system: Transaction Beginning Inventorc. Weighted Average Ending Inventory Cost of Goods Sold
Answer #1

Answer-

(a) FIFO

Ending inventory =$3,414

Cost of goods sold=$4,980

Working:

cost of goods sold = 16 x $66 +26 x $72+ 27x $76

=$4,980

Ending inventory= 5x $76+ 37x $82 =$3,414

(b) LIFO

Ending inventory=$3,172

Cost of goods sold=$5,222

Working:

Cost of goods sold= 26x$72+ 11x$66+ 32x $82

=$5,222

Ending inventory= 5x$66 + 32x$76 +5x$82 =$3,172

(C) weighted average

Ending inventory=$3,176

Cost of goods sold=$5,218

  

Working:

Cost of goods sold= 37x 75.62 + 32x 75.62= $5,218

Ending inventory= 42x 75.62= $3,176